OpenStudy (anonymous):

if 12 counters are 3/4 of a set what is the whole set

OpenStudy (anonymous):

12 divided by 3/4 = 12/(3/4) = 12 * 4/3 = 16

OpenStudy (lilg132):

12 / 0.75 = 16

jimthompson5910 (jim_thompson5910):

Let x = number of counters total So "if 12 counters are 3/4 of a set", then 3/4x = 12 Now solve for x 3/4x = 12 3x = 12*4 3x = 48 x = 48/3 x = 16 So there are 16 counters total

OpenStudy (anonymous):

thanks